# Multi-Level Damage-Aware Graph Learning for Resilient UAV Swarm Networks

### v2

@@ Abstract
Unmanned aerial vehicle (UAV) swarm networks leverage resilient algorithms to address communication network split issues and restore connectivity. However, existing graph learning-based resilient algorithms face over-aggregation and non-convergence problems caused by uneven and sparse topology under massive damage scenarios. To alleviate these problems, we propose a novel Multi-Level Damage-Aware Graph Learning (ML-DAGL) algorithm, which generates recovery trajectories by mining information from destroyed UAVs. We first introduce a Multi-Branch Damage Attention (MBDA) module, which forms a sequence of multi-hop Damage Attentive Graphs (mDAG) with different ranges of receptive fields. Each mDAG links only remaining and damaged nodes to ensure a more even degree distribution for mitigating over-aggregation, and utilizes multi-hop dilation to establish more links for sparse topology enhancement. To resort to the mDAG, we propose a Dilated Graph Convolution Network (DGCN), which generates the optimal recovery trajectories with theoretically proven convergence under massive damage cases. Simulation results show that the proposed algorithm can guarantee the connectivity restoration under large swarm and damage scales, while significantly expediting the recovery time by 75.4% and improving the topology uniformity after recovery.
Index Terms:
Resilient network, UAV swarm, deep learning, graph convolution network.
@@ I Introduction
Among existing optimal planning algorithms, graph-learning-based approaches [23, 24] have achieved state-of-the-art performance by considering the interactions between UAV nodes and leveraging the powerful information aggregation capabilities of Graph Convolution Networks (GCN). However, there are two major challenges in using graph-learning algorithms to address CNS issues. The first is spatial over-aggregation that arises from the uneven topology of the remaining network, where certain nodes become highly clustered. This can lead to some nodes focusing excessively on a few high-degree neighbors while neglecting others, resulting in inefficient recovery trajectories that aggregate around high-degree nodes. The second challenge is non-convergence, which occurs under massive damage scenarios where the network becomes sparse as few links remain within disconnected regions. This sparsity can impede the convergence of graph-based algorithms, as nodes lack sufficient local connectivity to effectively propagate information. Consequently, the recovery process may either fail to converge or take significantly longer, especially in the case of extensive damage.
In this paper, we propose a Multi-Level Damage-Aware Graph Learning (ML-DAGL) algorithm to overcome the challenges outlined above. With the help of satellite devices or ground stations, UAV nodes can easily obtain global information and leverage the network topology for local computing. The algorithm generates recovery trajectories from valuable information of destroyed UAVs through two steps. First, the Multi-Branch Damage Attention (MBDA) module generates a series of multi-hop Damage-Attentive Graphs (mDAGs), where each mDAG has different ranges of damage awareness. By selectively linking remaining and damaged nodes, mDAG ensures a more even degree distribution and helps mitigate spatial over-aggregation. The multi-hop dilation of mDAG also addresses sparse topologies by establishing links between distant nodes, improving receptive field enhancement in regions with severe damage. Second, the Dilated Graph Convolution Network (DGCN) utilizes these mDAGs to compute optimal recovery trajectories. By employing the Bipartite Graph Convolution Operation (B-GCO) and the residual connection mechanism, the DGCN can preserve the gradient and guarantee convergence even under fragmented topologies of extensive damage scenarios. Simulation results indicate that the proposed algorithm guarantees connectivity restoration across varying damage scenarios, significantly reduces average recovery time, and enhances the recovered topology with a more uniform degree distribution. Our contributions in this paper are summarized as follows.
@@
    We propose a novel ML-DAGL algorithm to address the CNS issue in the USNETs. The damage-attentive information can guide the remaining UAVs to fill up the damaged regions to mitigate over-aggregation, and the multi-hop dilation on damage awareness can establish more links to address the problem of sparse topologies.
@@
    To adapt to diverse damage cases with varying scales, we propose the MBDA module to form numerical mDAGs with different ranges of damage awareness via parallel dilation branches. This module provides multi-level receptive fields for incorporating both local and global network information.
@@
    To ensure connectivity restoration on sparse topology, we propose DGCN that utilizes the mDAGs to compute optimal recovery trajectories with theoretically proved convergence. This achieves more efficient and robust graph learning compared with using the original GCO in [25].
  4. 4.
    The proposed algorithm only needs 6.02 Mb parameters and 513 MiB running memory, which is feasible for deployment on practical UAVs. Moreover, the algorithm achieves state-of-the-art resilience based on global information, providing a reference baseline for research on the local-information guided recovery methods.
@@ II Related Work
Resilient recovery algorithms for USNET have become a critical area of research, particularly in scenarios where connectivity is disrupted due to massive node destruction in hostile environments [26]. Various strategies have been developed to address the CNS challenges, focusing on reconnecting divided sub-nets within minimum recovery time. Considering the different motivations for moving the surviving nodes, we can divide the existing mobility-based approaches into critical node replacement, central aggregation, and optimal planning algorithms.
@@ III-A UAV Communication Model
Assuming that the USNET is composed of $N$ massive independent and identical UAVs with a pre-defined connected structure, where the UAVs are labeled by a global index =$\mathcal{I}=\{1,2,...,N\}$. Generally, UAVs form a network at equal altitudes for most missions, hence we mainly consider two-dimensional locations. The location vector $\bm{p}_{i}(t)={[x_{i}(t),y_{i}(t)]}^{\top}\in\mathbb{R}^{2}$ denotes the coordinates of the $i$-th UAV at time $t$, where $x_{i}(t)$ and $y_{i}(t)$ denote the coordinate component of $X$ and $Y$ axis, respectively.
Since we focus on the CNS issue in USNETs, the transmission delay of the air-to-air communication channel is not considered in this paper. Without loss of generality, we consider a predominantly line-of-sight communication link between UAVs in the air. According to Friis transmission formula, the received signal power of $u_{j}$ from $u_{i}$ is given by
@@
where $P_{0}$ is the transmitting power, $G_{tx},G_{rx}$ are the constant gains of the transmitting and receiving antennas, respectively. The $L(d_{ij})=\left(\lambda_{c}/{4\pi d_{ij}}\right)^{2}$ is the large-scale fading, where $\lambda_{c}$ is the wavelength, $d_{ij}=\|\bm{p}_{i}-\bm{p}_{j}\|$ is the Eulerian distance between the UAV $u_{i}$ and $u_{j}$. The $|h_{0}|^{2}$ is the small-scale fading that follows the Gamma distribution. Denoting the receiver sensitivity $\tau$, we can calculate the maximum communication range from the following condition
$P_{0}G_{tx}G_{rx}L(d_{tr})|h_{0}|^{2}=\tau,$ (2)
and $u_{i}$ and $u_{j}$ can establish a communication link at time $t$ when the distance is shorter than the UAV transmission range $d_{tr}$, i.e., the edge $e_{ij}\in\mathcal{E}(t)$ exists if and only if $d_{ij}(t)\leq d_{tr}$.
@@ III-B Graph Topology of USNET
Let graph $\mathcal{G}(t)=\{\mathcal{U}(t),\mathcal{E}(t)\}$ denote the topology of the USNET, where $\mathcal{U}(t)=\{u_{i}|i\in\mathcal{I}(t)\}$ represents the node set and $\mathcal{E}(t)=\{e_{ij}|u_{i},u_{j}\in\mathcal{U}(t)\}$ represents the edge set of communication links. The edge $e_{ij}$ exists if and only if node $u_{i}$ and $u_{j}$ form a communication link. Generally, the USNET establishes an ad hoc network to enable message exchange across the multi-hop links. Denote $H_{ij}(t)$ as the number of hops between nodes $u_{i}$ and $u_{j}$ at time $t$, which is calculated as the minimum number of transmissions to reach each other. In this case, node $u_{j}$ is called a $k$-hop neighbor of $u_{i}$ at time $t$ if $H_{ij}(t)=k$.
The massive damage can occur at any time and destroy a certain number of UAVs. To simplify the damage model, we determine the moment $t_{0}$ that the communication network split (CNS) issue happens, where a total of $N_{D}$ UAVs were destroyed and $N_{R}=N-N_{D}$ UAVs remained. The damage index $\mathcal{I}_{D}=\{d_{1},d_{2},...,d_{N_{D}}\}$ labeled all destructed UAVs, and the remain index $\mathcal{I}_{R}=\{r_{i}|r_{i}\in\mathcal{I}-\mathcal{I}_{D}\}$ contained the remaining UAV nodes. To model the topology after the damage, we constructed a remaining graph $\mathcal{G}_{R}(t)=\{\mathcal{U}_{R}(t),\mathcal{E}_{R}(t)\}$ at time $t\geq t_{0}$, where $\mathcal{U}_{R}(t)=\{u_{r_{i}}|r_{i}\in\mathcal{I}_{R}\}$ is the remaining-node set, and $\mathcal{E}_{R}(t)=\{e_{r_{i}r_{j}}|e_{r_{i}r_{j}}\in\mathcal{E}(t),r_{i},r_{j}\in\mathcal{I}_{R}\}$ is the edge set for remaining UAVs.
In this paper, we assume that each UAV periodically reports its position and status information through satellite communication devices. Therefore, the satellite can sense the damaged UAVs and then broadcast the global information, including the damage index and position matrix. The position matrix consisting of the location vectors of each UAV at time $t$ is defined as
@@ III-C Network Connectivity
Due to the CNS issue, the remaining graph $\mathcal{G}_{R}(t)$ is split into several disconnected sub-nets. Denote $\mathcal{S}_{all}(t)=\{\mathcal{S}_{j}(t)|\mathcal{S}_{j}(t)\subseteq\mathcal{U}_{R}(t)\}$ as the series of sub-nets in $\mathcal{G}_{R}(t)$ at time $t$, where sets $\mathcal{S}_{i}(t)$ and $\mathcal{S}_{j}(t)$ are said to be disconnected at time $t$ if $\forall u_{k}\in\mathcal{S}_{i}(t)$ can not reach other nodes $\forall u_{l}\in\mathcal{S}_{j}(t)$.
Let $N_{s}(t)=|\mathcal{S}_{all}(t)|$ denote the number of sub-nets in $\mathcal{G}_{R}(t)$ at time $t$, apparently, $\mathcal{G}_{R}(t)$ is connected at time $t$ if and only if $N_{s}(t)=1$. A widely-used method to calculate $N_{s}(t)$ of $G_{R}(t)$ with the Laplace matrix is introduced in [28]. We first define the adjacency matrix of $\mathcal{G}(t)$ at time $t$ as $\bm{A}(t)=(a_{ij})\in\mathbb{S}^{N}$, where $a_{ij}$ is a 0-1 variable to denote if $e_{ij}$ exist or not, and $\mathbb{S}^{N}$ represents the set of symmetric matrices. Similarly, the adjacency matrix of the $\mathcal{G}_{R}(t)$ at time $t$ is defined as $\bm{A}_{R}(t)=(a_{r_{i}r_{j}})\in\mathbb{S}^{N_{R}}$, and $a_{r_{i}r_{j}}=1$ if and only if $e_{r_{i}r_{j}}\in\mathcal{E}_{R}(t)$. With the adjacency matrix $\bm{A}_{R}(t)$, we have the degree matrix of $\mathcal{G}_{R}(t)$ as
@@
where $d_{r_{i}}(t)=\sum_{j=r_{1}}^{r_{N_{R}}}a_{r_{i}r_{j}}(t)$ is the degree of each remaining node, and ${\rm diag}(\cdot)$ represents a diagonal matrix. The Laplace matrix of $\mathcal{G}_{R}(t)$ is define as the difference between $\bm{D}_{R}(t)$ and $\bm{A}_{R}(t)$, i.e.,
@@
Since the $\bm{L}_{R}(t)$ is a positive semi-definite matrix, with the eigenvalue decomposition applied on $\bm{L}_{R}(t)$, we have
@@
where $\bm{U}_{R}(t)=[\bm{u}_{r_{1}}(t),\bm{u}_{r_{2}}(t),...,\bm{u}_{r_{N_{R}}}(t)]$ as the unitary matrix composed of orthogonal eigenvectors, and $\bm{\Lambda}_{R}(t)={\rm diag}(\lambda_{r_{1}}(t),\lambda_{r_{2}}(t),...,\lambda_{r_{N_{R}}}(t))$ denotes the matrix with non-negative eigenvalues. The number of sub-net is equal to the number of zero eigenvalues in $\bm{\Lambda}_{R}(t)$, i.e.,
@@
Apparently, for the remaining graph $\mathcal{G}_{R}(t_{0})$ after the CNS issue, we have $N_{s}(t_{0})>1$.
@@ III-D Problem Formulation
Mobility-based resilient algorithms aim to restore the connectivity of $\mathcal{G}_{R}(t)$ with minimum recovery time via the UAV movement. For the original USNET with $N$ UAVs, there can be $2^{N}$ different damage cases under varying numbers of destroyed UAVs $N_{D}$. Note that not every case will split the network, we only study the damage scenarios in which CNS issues happen at time $t_{0}$. Denote $T_{r_{i}}$ as the flying time of the remaining node $u_{r_{i}}$ during the recovery process, we have the total recovery time as $T_{rc}={\rm max}_{r_{i}\in\mathcal{I}_{R}}T_{r_{i}}$, and the connectivity constraint $N_{S}(t_{0}+T_{rc})=1$ holds.
To find the recovery trajectories of remaining nodes, we need to generate an optimal target position matrix $\hat{\bm{X}}_{R}=[\hat{\bm{p}}_{r_{1}},\hat{\bm{p}}_{r_{2}},...,\hat{\bm{p}}_{r_{N_{R}}}]^{\top}$ for the target remaining graph $\mathcal{\hat{G}}_{R}$. The recovery process can be simplified as each node moves towards its target location straightly with its maximum speed $v_{max}$. Consequently, the flying time of the remaining node $u_{r_{i}}$ from its initial location to the target location is calculated by $T_{r_{i}}=\frac{1}{v_{max}}\|\bm{\hat{p}}_{r_{i}}-\bm{p}_{r_{i}}(t_{0})\|$. The total recovery time depends on the maximum time required for the remaining drones to reach their respective target positions. Thus, the problem can be reformulated as
[…]
@@
where $\hat{N}_{S}$ is the number of sub-nets in the target remaining graph $\hat{\mathcal{G}}_{R}$ recovered by $\bm{\hat{X}}_{R}$.

### v3

@@ Abstract
Unmanned aerial vehicle (UAV) swarm networks leverage resilient algorithms to restore connectivity from communication network split issues. However, existing graph learning-based approaches face over-aggregation and non-convergence problems caused by uneven and sparse topology under massive damage. In this paper, we propose a novel Multi-Level Damage-Aware (MLDA) Graph Learning algorithm to generate recovery solutions, explicitly utilizing information about destroyed nodes to guide the recovery process. The algorithm first employs a Multi-Branch Damage Attention (MBDA) module as a pre-processing step, focusing attention on the critical relationships between remaining nodes and destroyed nodes in the global topology. By expanding multi-hop neighbor receptive fields of nodes to those damaged areas, it effectively mitigating the initial sparsity and unevenness before graph learning commences. Second, a Dilated Graph Convolution Network (DGCN) is designed to perform convolution on the MBDA-processed bipartite graphs between remaining and destroyed nodes. The DGCN utilizes a specialized bipartite graph convolution operation to aggregate features and incorporates a residual-connected architecture to extend depth, directly generating the target locations for recovery. We theoretically proved the convergence of the proposed algorithm and the computational complexity is acceptable. Simulation results show that the proposed algorithm can guarantee the connectivity restoration with excellent scalability, while significantly expediting the recovery time and improving the topology uniformity after recovery.
@@ I Introduction
Among existing optimal planning algorithms, graph-learning-based approaches [32, 23, 24] have achieved state-of-the-art performance by considering the interactions between UAV nodes and leveraging the powerful information aggregation capabilities of Graph Convolution Networks (GCN). However, there are three major challenges in using graph-learning algorithms to address CNS issues. The first is spatial over-aggregation that arises from the uneven topology of the remaining network, where certain nodes become highly clustered. This can lead to some nodes focusing excessively on a few high-degree neighbors while neglecting others, resulting in inefficient recovery trajectories that aggregate around high-degree nodes. The second challenge is that a sparse network topology under extensive damage can impede the convergence of graph-based algorithms, as nodes lack sufficient local connectivity to effectively propagate information. Consequently, these algorithms may fail to restore the connectivity. The third challenge lies on the scalability. Since GCN is a transductive learning approach [25], the changes of damage cases and swarm scales require additional online iterations or numerous pre-trained models, harming the deployment feasibility of graph learning-based algorithms.
In this paper, we propose a Multi-Level Damage-Aware Graph Learning (ML-DAGL) algorithm that leverages location of destroyed UAVs to guide the recovery process, which generates recovery trajectories through two steps. First, a Multi-Branch Damage Attention (MBDA) module forms a series of multi-hop Damage-Attentive Graphs (mDAGs) that can overcome all three problems outlined above. By assigning the attention of surviving nodes to their destroyed neighbors, mDAGs form more even topologies and help mitigate spatial over-aggregation. The multi-hop dilation of mDAGs also addresses sparse topology issues by improving nodes’ receptive fields in severe damage regions. Moreover, the unified structure of mDAGs is insensitive to the scale of damage as they always considers all nodes, thus they have good scalability under damage-varying cases. Second, due to the special bipartite topology of mDAGs, we design a Dilated Graph Convolution Network (DGCN) to compute recovery trajectories based on mDAGs. The Bipartite Graph Convolution Operation (B-GCO) in DGCN efficiently aggregates features among remaining and destroyed nodes, and the residual connection mechanism preserves the gradient for deep structure. Theoretical analysis showed that the proposed algorithm has stable convergence and acceptable computational complexity. Simulation results indicate that the proposed algorithm guarantees connectivity restoration across varying damage scenarios, significantly reduces average recovery time, and enhances the recovered topology with a more uniform degree distribution. Our contributions in this paper are summarized as follows.
@@
    We propose a novel ML-DAGL algorithm for UAV swarm connectivity restoration, explicitly incorporating all UAV nodes, both surviving and destroyed, into a fixed-dimension representation as mDAG. This fundamental design decouples model input from specific damage patterns, enabling a single pre-trained model to handle any damage scale within a fixed swarm scale, directly solving the scalability challenge of GCNs.
@@
    We introduce the MBDA module to pre-process the global topology, with a novel damage attention mechanism that focus on critical surviving-destroyed node links. Combined with parallel multi-hop dilated processing, it actively reshapes sparse and uneven topologies into dense, uniform mDAGs before learning, inherently mitigating over-aggregation and information sparsity.
@@
    We design DGCN to operate on the mDAGs. It uses a bipartite graph convolution for feature aggregation from destroyed nodes to surviving nodes and a residual architecture for depth. Its theoretical convergence proof ensure stable training and reliable trajectory generation even under severe fragmentation, resolving the non-convergence challenge of prior GCNs.
@@ II Related Work
Resilient recovery algorithms for USNET have become a critical area of research, particularly in scenarios where connectivity is disrupted due to massive node destruction in hostile environments [26]. Various strategies have been developed to address the CNS challenges, focusing on reconnecting divided sub-nets within minimum recovery time. Considering the motivations for moving the surviving nodes, we can divide mobility-based approaches into critical node replacement, central aggregation, and optimal planning algorithms.
@@ III-A UAV Communication Model
Assuming that the USNET is composed of $N$ massive independent and identical UAVs with a pre-defined connected structure. Generally, UAVs form a network at equal altitudes for most missions, hence we mainly consider two-dimensional locations. The location vector $\bm{p}_{i}(t)={[x_{i}(t),y_{i}(t)]}^{\top}\in\mathbb{R}^{2}$ denotes the coordinates of the $i$-th UAV at time $t$, where $x_{i}(t)$ and $y_{i}(t)$ denote the coordinate component of $X$ and $Y$ axis, respectively.
Since we focus on the CNS issue in USNETs, the transmission delay of the air-to-air communication channel is not considered in this paper. Without loss of generality, we consider a predominantly line-of-sight communication link between UAVs in the air. According to Friis transmission formula, the received signal power of $u_{j}$ from $u_{i}$ is
@@
where $P_{0}$ is the transmitting power, $G_{tx},G_{rx}$ are the constant gains of the transmitting and receiving antennas, respectively. $L(d_{ij})=\left(\lambda_{c}/{4\pi d_{ij}}\right)^{2}$ is the large-scale fading, where $\lambda_{c}$ is the wavelength, $d_{ij}=\|\bm{p}_{i}-\bm{p}_{j}\|$ is the Eulerian distance between the UAV $u_{i}$ and $u_{j}$. The $|h_{0}|^{2}$ is the small-scale fading that follows the Gamma distribution. Denoting the receiver sensitivity $\tau$, condition of UAVs $u_{i}$ and $u_{j}$ that can form a communication link follows
$P_{0}G_{tx}G_{rx}L(d_{ij})|h_{0}|^{2}\leq\tau.$ (2)
We can then determine the maximum transmission range $d_{tr}$ of UAV by solving (2) as an equation. Therefore, the communication link between $u_{i}$ and $u_{j}$ exists if and only if $d_{ij}(t)\leq d_{tr}$.
@@ III-B Graph Topology of USNET
Let graph $\mathcal{G}(t)=\{\mathcal{U}(t),\mathcal{E}(t)\}$ denote the topology of the USNET, where $\mathcal{U}(t)=\{u_{i}|i=1,2,...,N\}$ represents the node set and $\mathcal{E}(t)=\{e_{ij}|u_{i},u_{j}\in\mathcal{U}(t)\}$ represents the edge set of communication links. The edge $e_{ij}$ exists if and only if node $u_{i}$ and $u_{j}$ form a communication link. Generally, the USNET establishes an ad hoc network to enable message exchange across the multi-hop links. Denoting $H_{ij}(t)$ as the number of hops between nodes $u_{i}$ and $u_{j}$ at time $t$, node $u_{j}$ is called a $k$-hop neighbor of $u_{i}$ at time $t$ if $H_{ij}(t)=k$.
With the help of satellite devices or ground stations, UAV nodes can easily obtain global information for local computing. In this paper, we assume that each UAV periodically reports its position and status information through satellite communication devices. Therefore, the satellite can record the node status as on-line and off-line, and summarize the global position matrix $\bm{X}(t)$ consisting of the location vectors of each UAV at time $t$, i.e.,
@@
The global information of node status and position matrix are then broadcasted to each UAV nodes periodically for local computation.
Notice that the damage can occur at any time during and destroy a certain number of UAVs, there can be more than $2^{N}$ different damage cases under varying numbers of destroyed UAVs. Since not every damage cases will split the network, we only study the damage scenarios in which can result in the communication network split (CNS) issues. The damage model in this paper is simplified as $N_{D}$ UAVs were destroyed at time $t_{0}$, and $N_{R}=N-N_{D}$ UAVs remained with CNS issues. Based on the status vector, we can label the damage index $\{d_{1},d_{2},...,d_{N_{D}}\}$ and the remain index $\{r_{1},r_{2},...,r_{N_{R}}\}$. To model the topology after the damage, we constructed a remaining graph $\mathcal{G}_{R}(t)=\{\mathcal{U}_{R}(t),\mathcal{E}_{R}(t)\}$ at time $t\geq t_{0}$, where $\mathcal{U}_{R}(t)=\{u_{r_{i}}|r_{i}=r_{1},r_{2},...,r_{N_{R}}\}$ is the remaining-node set, and $\mathcal{E}_{R}(t)=\{e_{r_{i}r_{j}}|e_{r_{i}r_{j}}\in\mathcal{E}(t),u_{r_{i}},u_{r_{j}}\in\mathcal{U}_{R}(t)\}$ is the edge set for remaining UAVs.
@@ III-C Network Connectivity
Due to the CNS issue, the remaining graph $\mathcal{G}_{R}(t)$ is split into several disconnected sub-nets. Let $N_{s}(t)$ denote the number of sub-nets in $\mathcal{G}_{R}(t)$ at time $t$, apparently, $\mathcal{G}_{R}(t)$ is connected at time $t$ if and only if $N_{s}(t)=1$.
A widely-used method to calculate $N_{s}(t)$ of $G_{R}(t)$ with the Laplace matrix is introduced in [28]. We first define the adjacent matrix of $\mathcal{G}(t)$ at time $t$ as $\bm{A}(t)=(a_{ij})\in\mathbb{S}^{N}$, where $a_{ij}$ is a 0-1 variable to denote if $e_{ij}$ exists or not, and $\mathbb{S}^{N}$ represents the set of $N\times N$ dimensional symmetric matrices. Similarly, the adjacent matrix of the $\mathcal{G}_{R}(t)$ at time $t$ is defined as $\bm{A}_{R}(t)=(a_{r_{i}r_{j}})\in\mathbb{S}^{N_{R}}$, and $a_{r_{i}r_{j}}=1$ if and only if $e_{r_{i}r_{j}}\in\mathcal{E}_{R}(t)$. With the adjacent matrix $\bm{A}_{R}(t)$, we can calculate degree matrix of $\mathcal{G}_{R}(t)$ as
@@
where $d_{r_{i}}(t)=\sum_{j=r_{1}}^{r_{N_{R}}}a_{r_{i}r_{j}}(t)$ is the degree of each remaining node, and ${\rm diag}(\cdot)$ represents a diagonal matrix.
The Laplace matrix of $\mathcal{G}_{R}(t)$ is then calculated as the difference between $\bm{D}_{R}(t)$ and $\bm{A}_{R}(t)$, i.e.,
@@
Note that $\bm{L}_{R}(t)$ is a positive semi-definite matrix, with the eigenvalue decomposition applied on $\bm{L}_{R}(t)$, we have
@@
where $\bm{U}_{R}(t)=[\bm{u}_{r_{1}}(t),\bm{u}_{r_{2}}(t),...,\bm{u}_{r_{N_{R}}}(t)]$ as the unitary matrix composed of orthogonal eigenvectors, and $\bm{\Lambda}_{R}(t)={\rm diag}(\lambda_{r_{1}}(t),\lambda_{r_{2}}(t),...,\lambda_{r_{N_{R}}}(t))$ denotes the matrix with non-negative eigenvalues. The number of sub-nets is equal to the number of zero eigenvalues in $\bm{\Lambda}_{R}(t)$, i.e.,
@@ III-D Problem Formulation
The goal of resilient algorithms is to restore the connectivity of $\mathcal{G}_{R}(t)$, i.e., decreasing $N_{S}(t)$ to 1. Compared with computing instant velocities for each UAVs during the recovery process, a simplified way is to generate a target position matrix $\hat{\bm{X}}_{R}=[\hat{\bm{p}}_{r_{1}},\hat{\bm{p}}_{r_{2}},...,\hat{\bm{p}}_{r_{N_{R}}}]^{\top}$, which forms a target connected graph $\mathcal{\hat{G}}_{R}$ with $\hat{N}_{S}=1$. Therefore, the recovery process is presented as each node moves towards its target location straightly with a maximum speed $v_{max}$ then hovers in place.
We now define the recovery time cost $T_{rc}$ to measure the efficiency of algorithms. We neglect inertial effects as it is often done in the control literature on swarm resilience [22, 23]. The flying time of the remaining node $u_{r_{i}}$ is given by
$T_{r_{i}}=\frac{1}{v_{max}}\|\bm{\hat{p}}_{r_{i}}-\bm{p}_{r_{i}}(t_{0})\|.$ (8)
Obviously, the total recovery time $T_{rc}$ depends on the maximum flying time of the final remaining UAV reaching its target position, i.e., $T_{rc}={\rm max}_{u_{r_{i}}\in\mathcal{U}_{R}}T_{r_{i}}$. Take $T_{rc}$ as the optimization target along with the connectivity constraint, the problem of CNS recovery is formulated as
$({\rm P}1):\quad\mathop{\rm min}\limits_{\bm{\hat{X}}_{R}}\quad$ $T_{rc}=\mathop{\rm max}\limits_{u_{r_{i}}\in\mathcal{U}_{R}}\frac{\|\bm{\hat{p}}_{r_{i}}-\bm{p}_{r_{i}}(t_{0})\|}{v_{max}}$ (9)